OpenStudy (anonymous):

What are the solutions of the system? y = x^2 - 2x -2 y = 4x + 5 A. (–1, 1) and (–7, –23) B. (–1, 1) and (7, 33) C. (–1, 33) and (7, 1) D. no solution I Thinks It's A! Is That Correct?

OpenStudy (anonymous):

substitute y=4x+5 into the longer equation to get: 4x+5=x^2-2x-2 combine like terms. x^2-6x-7=0 Factor. (x-7)(x+1)=0

sam (.sam.):

no , its not a

OpenStudy (anonymous):

oh :/ i sorta figured i did something wrong but my other guess was D. No solution

sam (.sam.):

there's a solution

OpenStudy (anonymous):

Is it (-1,1) and (7, 33)? :/

sam (.sam.):

yes

OpenStudy (anonymous):

Thanx (:

sam (.sam.):

welcome

OpenStudy (anonymous):

its B (-1,10 and(7,33)